OpenStudy (anonymous):

Write a two-step equation that needs the steps of addition then multiplication in order to be solved, and then solve it. Show your work.

OpenStudy (anonymous):

@Directrix @Easyaspi314 @JAMG @jigglypuff314

OpenStudy (anonymous):

Would 4x -5 = 11 work?

OpenStudy (anonymous):

NO BECUASE YOU WOULD NEED THE STEPS OF MULTIPLICATION SO MABEY 4/X-5=11

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

You mean (x/4) - 5 = 11. Perfect!

OpenStudy (anonymous):

Put parenthesis around the (x/4), like I did to show that it is x divided by 4.

OpenStudy (anonymous):

can you explain how you figure out the answer please, thanks! @Easyaspi314

OpenStudy (anonymous):

Add 4 to both sides...what do you get?

OpenStudy (anonymous):

well?

OpenStudy (anonymous):

wait what? explain how you get the answer. Its my last question and i need to go to bed! Please do it!

OpenStudy (anonymous):

Add 5 to both sides, you get (x/4) = 16 Now multiply both sides by 4, you get x = 64. (To undo divsion, you multiply; thats why I multiplied both sides by 4)

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

welcome.